OpenStudy (mackenzie2013):

Danielle has taken five math tests so far this year. The tests are out of twenty points, and she has gotten the following scores. 17, 19, 20, 14, 16 What must Danielle score on her sixth test in order to have a mean of 17.5? 17 18 19 20

OpenStudy (texaschic101):

(17 + 19 + 20 + 14 + 16 + x) / 6 = 17.5 (86 + x)/6 = 17.5 -- multiply entire equation by 6 86 + x = 105 x = 105 - 86 x = 19 Danielle would need a 19 on her 6th test to have a mean of 17.5